1 May 1968
GOLDEN DRAGONS conducted a CA with three companies to establish LZ Roberts. LZ was cold. Charlie Co lead elements while pushing southeast from LZ Roberts spotted 2-3 NVA on the trail. Results 1 NVA KIA. One steel pot CIA. They also found trail marking as follows: K-7-L. S2 reports it is L trail which leads to some higher Headquarters to the south. Alpha Co moving north from LZ Roberts. Alpha Co moving north from LZ Roberts found 2 bunkers and a tree hooch which looked to be a FA or OPO position. Approx 1915 hrs LZ Roberts received 26 incoming 82mm mortar rounds with neg casualties. Hq group, Rcn plt and Delta Co moved to Polei Kleng.
PLANS: Continue to clear and develop LZ Roberts. Charlie Co to establish company strong point vic 891929. Alpha Co to establish Company strong point vic 901935.
2 May 1968
GOLDEN DRAGONS continued to clear LZ Roberts. 4.2 mortar plat moved from Polei Kleng to LZ Roberts. Approx 1935 hrs one trip flare left on LZ SWINGER went off. Arty was called in with unknown results.
PLANS: Continue work on LZ Roberts. Hq group to move from Polei Kleng to LZ Roberts.
3 May 1968
GOLDEN DRAGONS continued to clear LZ Roberts. LTC BALL returned from RR and assumed command of the Bn. Due to a lack of aircraft the scheduled move from Polei Kleng to LZ Roberts was canceled.
PLANS: Move Hq group and Delta Co to LZ Roberts. Continue to improve fire base.
4 May 1968
Bn rear are rec'd incoming rds of all type. Results 11 US WIA'S.
GOLDEN DRAGONS continued to improve LZ Roberts. During the early morning hours the Bn rear located in KONTUM received an unknown number of mortar and rocket rounds. Results 11 US WIA.
5 May 1968
Damage to rear area Equip 4-3/4 t trks 2-1/4 t trk 1-2 1/2 t trk, 1 GP tent all was hit by WP, He type rds(60m, 82m, 122 rockets).
Sky spots 885946 6 may 0130, 883936 0150, 903904 0700, 913914 1630, 889899 1600, 855960 1800, 890882 1630. All infor passed on to all units.
GOLDEN DRAGONS continue to improve LZ Roberts. At approximately 0642 hrs, KONTUM came under rocket and mortar fire. Results were 4-3/4 ton trucks, 2-1/4 ton trucks, 1-2 1/2 ton trucks damaged, no US casualties. While moving E, Delta Co, spotted markings on a tree pointing off the trail. Checked the area with neg results. During the night, LZ Roberts received two incoming mortar rounds, from the SE. B/2/9 ret fire. Negative damage to LZ Roberts.
PLANS: Prepare to move to LZ Bunker Hill. Advanced party to dept and prepare site. Rcn moving SE to recon the approach to LZ Roberts.
6 May 1968
GOLDEN DRAGONS continued to prepare defensive positions on and near LZ Roberts. Delta Co while reconing to the W found one LAW along the trail. Law destroyed in place. BN advanced party moved to LZ Bunker Hill at approx 1300 hrs. Delta Co conducted combat asault vic 891928 OPCON to 1/12 Infantry.
PLANS: Return of advanced party from LZ Bunker Hill due to change in mission. Alpha Co to be lifted from Polei Kleng to LZ Roberts.
7 May 1968
C/1/14 at loc 891928 received one incoming rocket, B-40 type. Three (3) WIA with walking wounds. Dustoff requested at 0700 and completed at 0803 hrs. Fired at point blank range, the B-40 hit a log in front of a bunker, the 3 WIA were standing outside the bunker had overhead cover.
GOLDEN DRAGONS continued defensive operations around LZ Roberts. Early in the morning, Charlie Co received one incoming B-40 type rocket. Results 3 US WIA, unknown enemy casualties. BN received change of mission and was ordered to continue defense of LZ Roberts. Alpha Co returned OPCON to 1/14 at 1530 hrs.
PLANS: Continue local recon patrols around defensive positions of LZ Roberts.
8 May 1968
C/1/14 16 element contact with enemy at 897935. See wrap up.
B/1/14 ajax is 86. (strength)
Wrap up of C/1/14 contact: 16 element moved down trail to the N about 100 meters from last night lauger loc. 10 to 15 enemy tried to flank 16 element. Contact resulted in 5 NVA KIA dressed in green uniforms, steel pots, weapons. C/1/14 received 15 to 20 mortar rds estimated to be 82mm. Will turn results of AS and Arty as soon as area can be checked.
GOLDEN DRAGONS continued assigned mission around LZ Roberts. Charlie Co while conducting recon patrols made contact with approx 15 NVA KIA, neg US casualties. Arty and TAC Air were employed in support of Charlie Company.
PLANS: Continue to improve defensive positions around LZ Roberts.
9 May 1968
GOLDEN DRAGONS continued to defend LZ Roberts. Local patrols from 81, 82, and 83 elements had neg sighting. New MALARIA PREVENTION measures were instructed through the BN.
PLANS: Continue to improve LZ Roberts and defensive position IAW current directions from BDE.
10 May 1968
Intell rept indicates that enemy may use diversionary chemical attack against LZ Roberts, LZ Incoming, LZ Brillo Pad. Col Forrester desires to know status of Gas Masks on LZ's immediately. At 1955 hrs all units been issued masks.
Mask count turned in by D/1/14 was incorrect need 100 masks. Also be advised, will be conducting recon in 5 man groups in area -- BDE CO
GOLDEN DRAGONS continued defensive operations in assigned AO. Charlie Co conducted recon patrols to the S and W of their positions. During the night, Bravo Co had movement N of their position. Grenades were employed and movement stopped, unknown enemy losses.
PLANS: Bravo Co to sweep area where movement occurred. Charlie and Alpha Companies to continue local patrols.
11 May 1968
GOLDEN DRAGONS continued to recon and patrol area around their defensive position. During early morning hours, Alpha Co had one trip flare go off. There were neg sightings. After check in morning Alpha discovered that one claymore mine was missing. The local patrols had neg findings for the day.
PLANS: Continue to improve LZ Roberts and conduct any local recon patrols.
12 May 1968
GOLDEN DRAGONS continued to conduct recon patrols and improve LZ Roberts. Recon elements from Bravo Co found one steel pot (NVA) and approx 20 one man fighting positions at 900937. At approx 2030 hours LZ Roberts received 12 incoming 82mm mortar rounds on the NW side of the perimeter. Rounds fell short of the perimiter and there was negative damage.
PLANS: Continue to improve LZ Roberts and conduct local patrols around the LZ.
13 May 1968
GOLDEN DRAGONS continued operations in assigned AO. Elements of the 7/17 CAV working in the GOLDEN DRAGON AO received ground to air fire vic 822953. TAC Air was called in with unknown results. Local patrols around LZ Roberts reported neg sightings during the day.
PLANS: Conduct a CA with Charlie Co to reinforce 1/12 Inf at LZ Bunker Hill. Continue local patrols.
14 May 1968
GOLDEN DRAGONS continued to defend LZ Roberts. Charlie Co reported movement to the N of their position. Grenades were employed with unknown results. CA with Charlie Co was postponed due to lack of aircraft. Recon patrols reported neg sightings for the day.
PLANS: Conduct CA with Charlie Co to reinforce 1/12 Inf. Continue to recon area around LZ Roberts.
15 May 1968
GOLDEN DRAGONS continued to operate in assigned AO. Local patrols report neg findings for the day. Charlie Co while moving to LZ Roberts, spotted an NVA setting up a claymore type mine. NVA ran and mine was blown in place by Charlie element. Charlie Co returned to its old location to spend the night.
PLANS: Conduct a CA with Charlie Co from its present posititon to LZ Bunker Hill. Continue to send out local patrols.
16 May 1968
C/1/14 while opcon 1-12 has 4 WIA at LZ Brillo Pad indiv wounded by mortars. Evac at 0845. At 0915 has 1 more WIA slightly wounded by mortar fire. At 1200 they have 1 KIA & 1 WIA indiv were hit by R&R fire. Total 6 WIA & 1 KIA. 1 WIA 2/9.
C/1/14 changed frequency because enemy was jamming their call sign. Alternate freq is being used at this time.
GOLDEN DRAGONS continued to defend LZ Roberts. Early in the morning LZ Roberts received 12-15 rounds of 82mm on the NW portion of hill. Rds landed outside the perimeter, neg damage. Charlie Co CA to LZ Bunker Hill OPCON 1/12. Delta Co while OPCON to 1/12 had 1 KIA and 6 WIA due to mortar and BR fire.
PLANS: Continue to defend and patrol area around LZ Roberts.
17 May 1968
GOLDEN DRAGONS continued to defend LZ Roberts. Local patrols reported neg findings for the day. LZ Roberts received one incoming RR round. Round hit 50m short of bunker line. Neg damage.
PLANS: Continue to improve and defend LZ Roberts. Local recon patrols and OP to be put out by Alpha and Bravo elements.
18 May 1968
GOLDEN DRAGONS continue to operate in MacArthur AO. 7/17 Cav will be operating in our AO vic grid 8312, 8304, 9104, 9197 east to Highway 14 along to 7. Early in the morning A/1/14 LP's closed with negative results. A/1/14 placed OPO's out during the daylight hours. They thought they heard movement and spotted one enemy and took him under fire with negative results. Bet he didn't run completely out of sight. He waited to see if Alpha OP would chase him; it look as if he was leading them into an ambush. OP returned to LZ Roberts and called artillery on enemy location. B/1/14 OP has movement to their front at 1342 about 200m at an AZ of 60 degrees. C and D/1/14 OPCON to 1/12. A/1/14 during the night heard movement and saw a flashlight. They called in Arty and Hvy Mort. Negative lights. Movement continued on perimeter resulting in a 100% alert for all elements on LZ Roberts. A/1/14 LP had movement and blew their claymores and returned to fire base then called in Arty on suspected enemy location. At mid-night Alpha had more movement and called for illumination. Saw one enemy and took him under fire.
PLANS: Continue to send out small patrols. A/1/14 will check area where they had movement. B/1/14 will send OP to North.
19 May 1968
GOLDEN DRAGONS continue to improve defensive positions on LZ Roberts. At 0100 hours Alpha and Bravo reported movement on the perimeter. Alpha 26 called for Ill and opened up with small arms and threw 2 grenades. Alpha swept area in morning and found a blood trail and drag marks in area where claymore was exploded on the re-supply pad. They found a bag with 7 Claymore plungers and a AK-47 mag fully loaded. Alpha sent OP to the E and Bravo sent OP to the W.
PLANS: Continue to send LP out at night and OP's during the day. Alpha will operate to the N.
20 May 1968
GOLDEN DRAGONS continue to secure and defend LZ Roberts. During the early morning hours Alpha LP's to the S and W heard movement 50 meters outside of the perimeter. The LP to the S also spotted one enemy took him under fire and they will check area in the morning. Later on in the morning the OP 22 Alpha heard chopping about 500 meters to their front and they called in artillery. The 36 element had movement to the front about 75 meters and fired M-79 rounds.
PLANS: Alpha and Bravo will conduct local patrol in AO.
21 May 1968
GOLDEN DRAGON OP's continue to have movement at night around perimeter. 31 Alpha reprots movement to the W. Negative action taken at this time. At 0853 Alpha sent out two OP's. Bravo sent out three OP's early in the morning. Alpha 32 at 1120 found 4 or 5 butterfly bombs. Engineers will blow them in place. Negative action in AO the rest of the day.
PLANS: Continue to improve and secure LZ Roberts. Alpha and Bravo will continue to send out OP's.
22 May 1968
Rcn/1/14 while OPCON Task Force Bentz at loc 042986, 0945 hrs, had a small amount of enemy contact resulting in 1 VC KIA and 1 AK-47 CIA. VC was wearing green pants, black shirt and Ho Chi Minh sandals.
Tomorrow's move of D/1/14 will be conduct on 45.75 Admin Log for 1/12 Inf.
GOLDEN DRAGONS at 0137 saw Alpha 21 reporting movement on E side of position. They threw two grenades with neg results at this time. Alpha and Bravo will send out three OP's each. OP's closed without any incidents. At 2320 hrs Bravo reports movement on re-supply pad. Illumination was called in with negative movement or signs of movement.
PLANS: Prepare for move from LZ Roberts to LZ Bass.
23 May 1968
GOLDEN DRAGONS saw LP's close early in the morning without any reported incidents during the night. At 1215 Hq, Hvy mort, Commo, Med, Echo(-) conducted a secure move from LZ Roberts to LZ Bass. A total of 16 "Hooks" were required for the move. Completed at 1730 hours. Alpha conducted a Ext from LZ Roberts to LZ Bass the lift was completed at 1815 hours.
PLANS: GOLDEN DRAGONS will take over Task Force Bentz AO and Rcn/1/14 will revert back to our control and C/1/10 will become OPCON to 1/14 Inf. Alpha/1/14 will conduct a secure move to ZA 045972 and push out to the N. Bravo /1/14 will S&D to the south.
24 May 1968
GOLDEN DRAGONS assumed responsibility of Task Force Bentz AO at 0800 hrs. Bravo/1/14 is patrolling to the S. At 1029 C/1/10 found a mine while on a road sweep with the engineers. They will blow it in place. At 1310 hours A/1/14 was lifted from LZ Bass to VIC of ZA 045972 and they will move out to the N. At 2032 C/1/10 reports a trip flare going off and they received 6 rounds of small arms fire. Ill was called in with negative results. A/1/14 reports small light of perimeter, the light appears to be moving. They have thrown three grenades and will check area in the morning. Later in the evening at their night location they heard movement to the E side of the perimeter and fired M-79 rounds. Noise has subsided at this time. Rear area in KONTUM is receiving incoming rounds at 2310 hours. Negative details at this time.
PLANS: Continue to construct defensive positions at LZ Bass. C/1/10 will send one element to the N and one element to the S. A/1/14 will patrol to the N on the ridge line. B/1/14 will patrol to the south. They will have the C/1/10 element in his area working for hime. 10/C/1/10 escort for convoy.
25 May 1968
The 1/14at KONTUM received approx 30-81mm mortar rds the 1/14 suffered 3 KIA and 5 WIA.
A/1/14 at loc 037997 has received 9 mortar rounds (60mm) and sporadic M-79 fire from the NE. There are now believed to be between 5-8 enemy entrenched to ahead of Alpha. Gunships are taking heavy automatic weapon fire from ridge to the E. Believed to be platoon size or larger. Report 6 US WIA. Dustoff requested at 0910 hrs completed at 1413. Dustoff 30 took rds and had to be relieved by Dustoff 37.
D/1/14 while OPCON 1/12 had 1 US WIA by small arm puncture. Not serious.
Alpha 6/1/14 reports a total of 9 WIA. He has a very slight wound from a piece of shrapnel from a M-79 round. It was just a slight wound and all it required was a band aid.
Rcn/1/14 found 2 dead NVA in a grave, dressed in green fatigues and wearing Ho Chi Minh sandals. Graves were about 5 weeks old. Graves were at location 038009.
10/C/1/10 at 048928 while on convoy found 3 dead Vietnamese - 1 male, 1 female, 1 child shot by small arms sometime after 1200 hrs.
A/1/14 receiving small arms fire from one enemy in spider hole. 36/A/1/14 moving in on him at this time. Action resulted in 2 US WIA - one serious (chest wounds) the other ambulatory. Dustoff requested.
A/1/14 reports that the WIA with chest wounds has become KIA. 36/A/1/14 has 4 enemy KIA and still is getting resistance from finger to the E. Receiving M-79, small arms and 2 mortar rounds at 037998/
A/1/14 at loc 037000 received incoming mortar rounds 2 US KIA
GOLDEN DRAGONS at 0045 hours saw 1/14 area at KONTUM receiving 30 rounds of 81mm mortar and they suffered 3 KIA and 5 WIA. The 20/C/1/10 is moving to the S and 30/C/1/10 is moving to the N at 0730. 81/A/1/14 is moving to the N patrolling the ridge line while 82/A/1/14 is moving to the south west. 26/A/1/14 reports receiving 10 rounds of AK-47 fire from their N. At 0710 A/1/14 is in contact with 5-8 enemy and they are receiving 60mm mortar rounds and sporadic M-79 fire. Gunships were called and arrived at the contact area while they were making their passes they received heavy automatic weapons fire. Size of enemy has increased and it is believed that the force is platoon size or larger. During this contact, A/1/14 received 8 US WIA. Dustoff was requested at 0910 and couldn't be complete until 1413 because of intense automatic and small arms fire on the LZ. Dustoff 30 tried to get into LZ but he took 6 rounds and had to return to Polei Kleng. The pilot requested another ship and completed the dustoff at 1413 hours. At 1130 A/1/14 can not maneuver because of enemy situation they are waiting for gunships. Rcn/1/14 made a CA behind the enemy, at 1155 hours. This maneuver took some of the pressure off A/1/14. Alpha 6 requested a new radio and RTO for his FO. At 1212, the Hvy Mort Plat was lifted from LZ Bass to LZ in vac ZA 046973 to provide support for A.1.14. Rcn/1/14 found 2 dead Vietnamese. At 1643 A/1/14 received sniper fire from an enemy in spider hole. While Alpha 36 was momving on enemy they received two US WIA one serious with a chest wound. Dustoff called at this time. The WIA with chest wounds is reported as KIA. Dustoff made an attempt to evacuate US WIA's. He received ground to air fire and was unable to complete evacuation. A/1/14 again came under enemy mortar fire resulting in two US KIA and one WIA. At 1530 hours 10/C/1/10 saw a VC running and took him under fire. Checked area with negative results. LZ Bass received 12 122mm rockets at 1822 negative casualties. Artillery was called on enemy with one secondary explosion reported.
26 May 1968
D/1/14 has 2 US KIA and 3 WIA from yesterday Ind were evacuated by 1-12 Inf.
A/1/14 found two more NVA KIA from yesterday's action.
The EM that was sent from D/2/35 to A/1/14 position yesterday was a WIA. Ind was extracted by C&C. Ind did an outstanding job for A/1/14.
Dustoff are extracting US WIA from A/1/14 that were wounded in yesterday action. Comp at 0840 hrs. Also 2 gunships escorted them.
A/1/14 found at loc 038998 1 NVA KIA. Ind was dressed in green also 1 Chicom grenade & 1 MG.
A/1/14 element found at loc 039998 1 NVA KIA ind was dressed in green w/1 chicom grenade.
26, A/1/14 at loc 038998 found a local national w/a PRC 25 on the LRRP Freq.
Cpt. Sells went to A/1/14 loc to pick up 4 US WIA ,3 Wia from yesterday and 1 today.
GOLDEN DRAGONS early in the morning 2C/1/10 reported seeing lights coming from the village. Fired Ill with neg results. 0622 LP's closed back to LZ Bass. 30C/1/10 has the mission of sweeping the road with a squad of engineers. A/1/14 found bodies of the enemy while sweeping yesterdays contact area. They also found one AK-47 and a light machine gun. CC went into A/1/14 location to pick up WIA's from yesterday contact. While the CC was going into A/1/14 location a man ran into the rotar blade and was KIA. D/1/14 has 2 KIA and 3 WIA from 25 May while OPCON 1/12 Inf. D/2/35 sent an EM to act as a guide for A/1/14 while he was doing his job he was WIA and evacuated. At 1040 hours a fake CA took place along with the insertion of the LRRP's. LRRP element will assume B/1/14 call sign and freq. PRU's (local forces) will operate in our AO from 26/1600 hrs to 31/1200.
PLANS: Continue to secure LZ Bass. A/1/14 will send patrols to the N and NE. B/1/14 will send patrols to the S. C/1/10 will send one element to the south and one element to the N. Rcn/1/4 will conduct local patrols around fire base. 1/14 will continue to send out OP and LP's.
27 May 1968
36/A/1/14 at loc 037904 found a spider hole booby trapped with a Chi Com grenade, 3 entrenching tools, a pack, and some documents also 1 dead NVA.
GOLDEN DRAGONS saw LP's close at 0600 hours neg activity during the night. 20C/1/10 is moving to the S. 26A/1/14 moving to the E. 10C/1/10 moving to the N. While 36A/1/14 was to the NE, they found a spider hole with a booby trapped Chicom grenade, 3 entrenching tools, a pack, some documents, and one NVA body. 26A/1/14 found a bunker complex at ZA 043997 with one enemy pack. The complex isn't new but it shows signs of recent use. 14 LRRP reported that they received 10 to 15 rounds vic ZA 012934. They moved to 30C/1/10 location and moved with 30C to sweep area with neg results. 26 and 36A/1/14 have found a new bunker complex vic ZA 047998 consisting of 50 bunkers with the capability of holding 750 enemy. CPT CARTER requested engineers to blow the complex.
PLANS: Continue to secure LZ Bass. A/1/14 will send patrols to the E and W. B/1/14 will send patrols SW and N. C/1/10 will conduct mine sweep and will send out OP's and LP's. Rcn/1/14 will move to the W.
28 May 1968
GOLDEN DRAGONS early in the morning saw A/1/14 report movement on south side of perimeter. Threw a couple of grenades with neg results. 0200 40/C/1/10 is receiving AW fire. They returned fire and enemy fire ceased. 10/C/1/10 has movement to the NE; they called Ill with neg results. 0640 30 and 10/C/1/10 moved out on mine sweep along with engineers. B/1/14 sends the 10 element to SW and 20 to the N. A/1/14 sent 16 to the N and 36 to the W. 0930 hrs Rcn/1/14 vic ZA 033967 found enemy camp site with what looks to be a mine field of "Bouncing Betty" mines along with some trails running N and S. They are calling Hvy Mort in on the area. At 1225 A/1/10 arrived at LZ Bass at 1410. LRRP(Bde) 14 element closed LZ Bass at 1425, and then moved out to vic ZA 013925. 2200 10/C/1/10 heard movement and called in Ill with neg results. Later at night 10/C/1/10 spotted light near water point. 2320 30/C/1/10 had movement to his front called in Hvy Mort with neg results.
PLANS: Continue to secure LZ Bass. A/1/10 will send out patrols ot N and W. They will put out LP's at night. C/1/10 will take care of mine sweep along with a team of engineers. They will also send two platoons to assist 2/1/14 in the move to LZ Brillo Pad. Rcn/1/14 will move to the S and then return to night location. A/1/14 will take engieers to bunker complex to destroy bunkers.
29 May 1968
GOLDEN DRAGONS saw B/1/14 move to the N to relieve D/1/14 at LZ Brillo Pad. 20 and 30/C/1/10 are moving to the S to assist B/1/14. 60/A/1/10 is moving to the N. Rcn/1/14 is moving S and then they will return to night location. 1/A/1/10 will move to the W. The observation plane spotted 35 people vic ZA 996883 mostly women and children. 1217 hrs A/1/14 along with engineers are moving to blow bunkers that were found yesterday. They blew the bunkers and returned to night locaion while engineers returned to LZ Bass. LRRP (Bde) element returned LZ Bass 20 AND 30/C/1/10 closed LZ Bass. B/1/14 unable to relieve D/1/14 will continue tomorrow.
PLANS: Continue to send OP's and LP's to the W and NW of LZ Bass. C/1/10 will sweep road along with engineers and they will send 2 elements to help B/1/14 make the switch with D/1/14. B/1/14 will try to relieve D/1/14. A/1/14 will send out patrols to the N and S. Rcn/1/14 will conduct local patrols around night loc.
30 May 1968
D/1/14 while OPCON 1/12 Inf has one US WIA named Dial. He has a slight shrapnel wound. Evacuated 1000 hrs 30 May.
GOLDEN DRAGONS: 61(-)/A/1/10 has to go back to night LP to retieve one track. 15/C/1/10 and the engineers are sweeping the road at 0805. 20 and 30 C/1/10 are moving south to assist in the change of B/1/14 with D/1/14 at LZ Brillo Pad. Today at 0835 1/14 received an extention. 26/A/1/14 is moving to the S, while 36/A/1/14 is moving to the W. Mine sweep has reached KONTUM at 1025. D/1/14 while OPCON 1/12 Inf received one WIA. 60/A/1/10 had a man fall off an APC. He is having difficulty in breathing. Dustoff requested and complete at 1110 hrs. D/1/14 OPCON 1/12 Inf while sweeping around LZ Brillo Pad found 1 AK-47, and 1 B-40 Rocket Launcher. Rcn has moved to the N while patrolling they found some dud firecracker rounds. At 1230 B/1/14 became OPCON 1/12 Inf. LRRP AO is 9894, 9892, 0092, and 0094. 26A/1/14 had neg finding in area where an AS had taken place. 26A/1/14 and 36A/1/14 have closed night location.
PLANS: Continue to secure LZ Bass. 61A/1/10 will put out local LP's. A/1/10 will conduct patrols to the N. C/1/10 will send elements to SW to assist in B/1/14 change with D/1/14. A/1/14 will send patrols to NW and SW. B/1/14 will move to LZ Brillo Pad to relieve D/1/14.
31 May 1968
GOLDEN DRAGONS today at 0637, 61/A/1/10 LP's closed LZ Bass with negative activity during the night. 20 and 30C/1/10 have departed to SW to assist B/1/14 and D/1/14 in switch at LZ Brillo Pad. LRRP found well used trail and followed it with neg results. 60A/1/10 moving to the N, 16 and 36A/1/14 moving to NW and SW respectively. 7/17 Cav AO 2200, 2208, 0608, 0503, 1700, 2210. 16A/1/14 smelled smoke and called in Hvy Mort with neg results. 20 and 30C/1/10 conducted local patrols while the wait for D/1/14. D/1/14 reverts to 1/14 Inf control and links up with 20 and 30/C/1/10.
PLANS: Continue to secure LZ Bass. C/1/10 along with engineers will conduct mine sweep. A/1/10 will conduct patrols to the N. A/1/14 will conduct patrols to the W. D/1/14 will conduct CA. A/1/10 will continue to put out LP's.